OpenStudy (anonymous):

Tina's team entered a math contest where teams of students compete by answering questions that are worth either 3 or 5 points. No partial credit is given. Tina's team scored 44 points on 12 questions. How many 5-point questions did the team answer correctly?

OpenStudy (anonymous):

7(5)+3(3)=44

OpenStudy (whpalmer4):

@Luis_Rivera let f be the number of 5 point questions, and t be the number of 3 point questions \(f + t = 12\) (12 questions altogether) \(5f + 3t = 44\) (44 points scored) \[f = 12 - t\]\[5(12-t) + 3t = 44\]\[60 -5t + 3t = 44\]\[-2t + 16 = 0\]\[t = 8\]\[f = 12 - 8 = 4\] So they answered 8 3 point questions and 4 5 point questions.

OpenStudy (whpalmer4):

Always check your work against the problem. Luis' solution is incorrect because it doesn't involve 12 questions, it involves 7+3 = 10 questions. 8+4 = 12, 8*3 + 4*5 = 24 + 20 = 44.